I have three questions if that is ok.

Solve for x: 5x - 12 = 4x + 5
Solve for x: -3(4x+3)-2 = 7 -2x + 3
Solve for x: (x + 3)/4 - 3/2 = (x + 4)/8
thank you in advance for you help

Answer by checkley77(12844)

I'LL LEAVE THE PROOFS TO YOU: IT'S GREAT PRACTICE.
------------------------------------------------
5X-12=4X=5
5X-4X=5+12
X=17 ANSWER.
--------------------
-3(4X+3)-2=7-2X+3
-12X-9-2=-2X+10
-12X+2X=10+11
-10X=21
X=21/-10 ANSWER.
-----------------------
(X+3)/4-3/2=(X+4)/8
(X+3)/4-(X+4)/8=3/2
[2(X+3)-(X+4)]/8=3/2
[2X+6-X-4]/8=3/2
[X+2]/8=3/2 NOW CROSS MULTIPLY.
2[X+2]=8*3
2X+4=24
2X=24-4
2X=20
X=20/2
X=10 ANSWER.